what is the solution of a triangular pennant is to be made with a base of 3 dm and a height of 4dm. It is to be cut from a square piece of cloth with a side of 4 dm, what is the area of a pennant? of the square piece of cloth? how much material will be left?
1
Expert's answer
2020-01-21T09:35:28-0500

SLKC=1/2LC*DC

SLKC=1/2*3*4=6dm2 (pennant area)

SABCD=(DC)2

SABCD=42=16dm2 (square area a piece of cloth)

SABLK+SCKD= 16dm2-6dm2=10dm2 (material which will be left after cutting).